### v2

# Ruiyan Han2∗  Zhen Fang1  Xinyu Sun2∗  Yuchen Ma2  Ziheng Wang2  Yu Zeng122footnotemark: 2 Zehui Chen1   Lin Chen1   Wenxuan Huang3,4   Weijie Xu5   Yi Cao6   Feng Zhao1 1

Abstract
While Unified Multimodal Models (UMMs) have achieved remarkable success in cross-modal comprehension, a significant gap persists in their ability to leverage such internal knowledge for high-quality generation. We formalize this discrepancy as Conduction Aphasia, a phenomenon where models accurately interpret multimodal inputs but struggle to translate that understanding into faithful and controllable synthesis. To address this, we propose UniCorn, a simple yet elegant self-improvement framework that eliminates the need for external data or teacher supervision. By partitioning a single UMM into three collaborative roles: Proposer, Solver, and Judge, UniCorn generates high-quality interactions via self-play and employs cognitive pattern reconstruction to distill latent understanding into explicit generative signals. To validate the restoration of multimodal coherence, we introduce UniCycle, a cycle-consistency benchmark based on a $Text\rightarrow Image\rightarrow Text$ reconstruction loop. Extensive experiments demonstrate that UniCorn achieves comprehensive and substantial improvements over the base model across six general image generation benchmarks. Notably, it achieves state-of-the-art (SOTA) performance on TIIF(73.8), DPG(86.8), CompBench(88.5), and UniCycle(46.5), while further delivering substantial gains of +5.0 on WISE and +6.5 on OneIG. These results highlight that our method significantly enhances T2I generation while maintaining robust comprehension, demonstrating the scalability of fully self-supervised refinement for unified multimodal intelligence.
[Uncaptioned image] UniCorn: Towards Self-Improving Unified Multimodal Models through Self-Generated Supervision
[…]
1 Introduction
The realization of Artificial General Intelligence (AGI) requires a tight synergy between comprehension and generation, wherein comprehension enables the internalization of knowledge and generation allows its coherent and expressive externalization. By integrating multiple modalities into a shared representational space, Unified Multimodal Models (UMMs) Deng et al. (2025b); Chen et al. (2025b); Xie et al. (2025c) naturally couple comprehension and generation as two complementary phases of a unified cognitive process, supporting both knowledge grounding and coherent reasoning.
Despite these advances, a fundamental disparity remains between comprehension and generation in current UMMs. This mismatch, which we formalize as Conduction Aphasia, arises when a model demonstrates strong domain understanding yet fails to translate that knowledge into high-quality generative outputs. As shown in Fig. 1, a representative case appears in image generation: although the model can accurately recognize what an image depicts and reliably assess its visual quality, it often cannot act on this knowledge during generation. This disconnect motivates a central research question: how can a model’s robust understanding guide and strengthen its generative behavior?
Driven by this simple yet fundamental question, we propose UniCorn, a post-training framework that enables self-improvement through a unified cycle of proposal, execution, and evaluation. Requiring no external data or teacher-model supervision, UniCorn allows UMMs to autonomously narrow the comprehension–generation gap by acting as their own instructor within a single parameter space. Motivated by the observation that a single UMM can exhibit distinct capabilities for proposing, executing, and evaluating, we treat the model as a modular system in which comprehension can explicitly guide generation. This design turns the model’s latent interpretive capability into an internal training signal, enabling autonomous generative improvement without external supervision.
Specifically, UniCorn operates through a self multi-agent framework that functionalizes the UMM into three distinct internal roles. The process begins with the model acting as a Proposer to propose diverse and expansive prompts, followed by its transition into a Solver to synthesize corresponding image candidates. Finally, it assumes the role of a Judge to provide evaluative rewards based on its superior comprehension.
Figure 1: Motivation of UniCorn. UMMs often exhibit an understanding-generation gap: they can accurately understand and critique errors in an image, yet fail to generate the same scene correctly. This conduction aphasia motivates our framework to leverage the model’s superior internal understanding to strengthen and refine its generative capabilities through self-contained feedback.
By simulating structured collaboration within a single parameter space, this design yields rich interaction data that are refined through data reconstruction. Concretely, we convert raw multi-agent outputs into structured training signals, including descriptive captions, evaluative judgments, and reflective feedback, thereby distilling latent understanding into explicit supervision for effective self-improvement.
To determine whether internal collaboration produces general multimodal intelligence instead of narrow task fitting, we introduce UniCycle, a cycle-consistency benchmark that probes cognitive alignment via informational integrity. Existing evaluations often separate comprehension and generation, which can lead to piecemeal measurements and biased conclusions. In contrast, UniCycle frames evaluation as a Text $\rightarrow$ Image $\rightarrow$ Text reconstruction process. It compares the model’s original intent with its reconstructed description, using the resulting semantic gap as a holistic, training-free indicator of conceptual coherence, while reducing the bias that arises when capabilities are tested in isolation.
Across extensive experiments, we find that our model achieves reliable self-improvement without heuristic reward engineering, curriculum design, or external supervision. Compared with prior self-improvement approaches Jin et al. (2025a) and methods that depend on external guidance, our approach learns from internally generated training signals, generalizes well, and remains stable under out-of-distribution (OOD) conditions. These results support the effectiveness of a fully self-contained learning paradigm.
Figure 2: Visualization results of UniCorn.
  • We identify the Conduction Aphasia phenomenon in UMMs, where strong understanding fails to translate into accurate generation, and propose UniCorn, which repurposes internal comprehension as self-supervision through Proposer, Solver, and Judge roles with data reconstruction.
  • To assess whether multimodal understanding and generation remain conceptually consistent across modality transitions, we introduce UniCycle, a training-free evaluation protocol that measures multimodal coherence through a Text $\rightarrow$ Image $\rightarrow$ Text cycle.
  • Experimental results demonstrate that our method consistently outperforms prior approaches, achieving SOTA performance on TIIF (73.8), DPG (86.8), CompBench (88.5), and UniCycle(46.5), together with substantial improvements of +4.0 on Geneval, +5.0 on WISE, and +6.5 on OneIG.
2 Related Work
Unified Multimodal Models
UMMs aim to unify cross-modal understanding and generation, yet strong understanding often fails to yield equally strong native generation. Existing designs fall into two paradigms: pure autoregressive models that jointly predict text and visual tokens over interleaved sequences  Chen et al. (2025e); Cui et al. (2025); Tong et al. (2025)) and hybrid models that combine autoregressive language modeling with diffusion-based image synthesis, either within a unified backbone  Xie et al. (2024); Zhao et al. (2024)) or via modular routing and sparse experts Shi et al. (2024); Liang et al. (2024b); Deng et al. (2025a)), with related guidance schemes such as Diffusion Forcing Chen et al. (2024a). Beyond architecture, self-improvement methods convert self-generated signals into training objectives Yu et al. (2025); Zhou et al. (2024b); Wang et al. (2025b); for UMMs, SRUM derives internal rewards from understanding Jin et al. (2025a), and UniRL jointly optimizes understanding and generation Mao et al. (2025). However, most pipelines depend on auxiliary components or task-specific feedback, limiting scalability and generalization.
Multi-Agent and Self-Improvement Learning
Multi-agent systems decompose reasoning through role specialization and interaction, enabling solution diversity and cross-verification, but often incur high coordination cost and brittle verification Chen et al. (2024d); Liang et al. (2024a); Cemri et al. (2025). In parallel, LLM self-improvement converts self-generated tasks and evaluations into training signals, supporting zero-data learning via self-play and self-rewarding mechanisms Silver et al. (2017); Huang et al. (2025a); Zhao et al. (2025a); Yuan et al. (2024). Unified Multimodal Models (UMMs) naturally unify understanding and generation within a single parameter space, making them particularly well-suited for lightweight role instantiation and fully model-driven self-improvement without external supervision.
3 Method
In this section, we begin by presenting the motivation through an analysis of the mismatch between generation and understanding capabilities in UMMs. Building on these observations, we introduce UniCorn, a simple yet elegant post-training framework that enables self-improvement without any external annotated data or teacher models.
3.1 Motivation
Just as a child who associates the word “apple” with the fruit can spontaneously name it upon seeing it, cognitive symmetry Blanco (2018) enables a bidirectional mapping between internal concepts and external expressions. This alignment is reminiscent of escaping Plato’s Cave: true intelligence must move beyond observing surface data to mastering the reciprocal relationship between an appearance and its underlying source.
However, current UMMs suffer from a functional deficit akin to Conduction Aphasia: while the model exhibits profound comprehension, its generative performance remains fractured, failing to produce the very content it can inherently understand. Bridging this gap is critical; without aligning these dual processes, a model remains a "passive observer," capable of grounding symbols but incapable of utilizing them. Mastering the synergy between understanding and generation is thus not merely a functional upgrade but the essential step toward achieving the cognitive integrity required for AGI.
On the one hand, as illustrated in Fig. 3, current UMMs demonstrate formidable perception and comprehension capabilities. Specifically, when serving as a reward model for Text-to-Image (T2I) generation, the UMM exhibits a sophisticated grasp of cross-modal semantics. This suggests that the model has already internalized a robust ’world model’ and possesses the necessary latent knowledge to discern high-quality visual-textual alignments.
Figure 3: Results of BAGEL Deng et al. (2025b) and GPT-4o Hurst et al. (2024) on four understanding benchmarks. For Omini-RewardBench Jin et al. (2025b) and MMRB2 Hu et al. (2025), we evaluate the T2I task. Performances are normalized with GPT-4 Achiam et al. (2023) results for better visualization.
On the other hand, the model’s generative capability remains markedly constrained, primarily due to its failure to bridge the gap between internal recognition and active synthesis. This functional dissociation means that the UMM’s own sophisticated understanding remains a ’silent passenger’ during the generative process, unable to inform or correct its outputs. Building on this observation, our key insight is that the UMM’s formidable comprehension can be repurposed as an autonomous supervisory signal to steer its generative behavior. By transforming latent interpretive depth into explicit guidance, we promote a tighter coupling between these two processes, ultimately restoring the cognitive symmetry essential for a truly integrated multimodal intelligence.
3.2 Problem Definition
Figure 4: Overview of the UniCorn Framework. (a) Illustrates the self-multi-agent collaboration for high-quality data sampling. (b) Details the Cognitive Pattern Reconstruction process, which reorganizes data to facilitate robust and efficient learning. (c) Presents the UniCycle benchmark evaluation, verifying whether the model can accurately reconstruct key textual information from its own generated content.
We study UMMs that process interleaved image-text inputs and outputs. A UMM is formulated as a policy $\pi_{\theta}$ that maps a multimodal input sequence
$X=(x_{1},\ldots,x_{N}),\quad x_{n}\in T\cup I,$ (1)
to an interleaved multimodal output sequence $Y=\pi_{\theta}(X)$. This unified input-output formulation supports both Image-to-Text (I2T) understanding and Text-to-Image (T2I) generation. We operationalize understanding as I2T and generation as T2I, and leverage the model’s stronger I2T understanding to supervise and refine its weaker T2I generation.
3.3 UniCorn
UniCorn operates via two core stages: Self Multi-Agent Sampling and Cognitive Pattern Reconstruction (CPR). First, the UMM concurrently assumes three roles: Proposer, Solver, and Judge (§ 3.3.1), to simulate a collaborative loop. Then, the CPR stage reconstructs these raw interactions into three training patterns: caption, judgement, and reflection (§ 3.3.2), which are combined with high-quality self-sampled T2I generation data for post-training. Critically, the entire process is fully self-contained, requiring no external teacher models or human-annotated data.
3.3.1 Stage 1: Self Multi-Agent Sampling
LLMs are naturally suited for self-play in multitask settings Radford et al. (2019). For UMMs, interleaved multimodal inputs and functional diversity allow prompting, generation, and judgement to coexist within a shared model, enabling role-conditioned behaviors under different prompts. We leverage this property to functionalize a single UMM into collaborative roles, bridging the comprehension–generation gap through internal synergy.
Proposer $\pi_{\theta}(T\mid T)$
The proposer is designed to generate a diverse set of challenging prompts for the unified multimodal model, which are subsequently used to produce training images. To this end, inspired by LAION-5B Schuhmann et al. (2022) and COYO-700M Byeon et al. (2022), we partition all T2I task prompts into ten categories and designed fine-grained generation rules for each category. Next, we prompt UMM to generate an initial batch of prompts and act as the judge to select the best candidate for subsequent iterations. Leveraging the strong in-context learning (ICL) capabilities of LLMs Dong et al. (2024), the initial example serves as a few-shot demonstration to guide the generation of subsequent prompts. To further enhance diversity, we introduce a dynamic seeding mechanism. After generating a predefined number of prompts, several examples are sampled from the prompt library for evaluation and then used to construct new demonstrations that guide the next round of prompt generation. Compared with prior approaches that either directly rely on training set Jin et al. (2025a) or employ external models for prompt construction Mao et al. (2025), our method requires no external data and generates more diverse prompts, thereby improving generalization.
Solver $\pi_{\theta}(I\mid T)$
The solver is responsible for producing a diverse set of outputs in response to the prompts generated by the proposer. Therefore, we encourage the UMM to generate images under random seeds and different hyperparameters. Following DeepSeek-R1 Guo et al. (2025a), we perform 8 rollouts per prompt to strike a favorable trade-off between sample quality, diversity, and computational efficiency.
Judge $\pi_{\theta}(T\mid T,I)$
The judge is responsible for assigning scores to the images generated by the solver in response to prompts proposed by the proposer, which are then used for rejection sampling during training.
Previous work has relied on heuristic reward functions based on keywords Mao et al. (2025) or on powerful external models to provide dense reward maps Jin et al. (2025a). Such reward judges depend heavily on parameter tuning and the performance of external models, which varies across tasks, thereby severely limiting the generalization of self-improvement. As illustrated in Fig. 3, UMMs exhibit strong reward modeling capabilities. Thus, we formulate reward evaluation for all T2I tasks using discrete scores ranging from 0 to 10, following a widely adopted LLM-as-a-judge paradigm Radford et al. (2019); Kim et al. (2023). To further enhance judgement quality, we transfer generation reward models Liu et al. (2025), which have demonstrated strong potential in LLMs, to T2I evaluation. Specifically, we design task-specific rubrics for each category and encourage the model to explicitly articulate its reasoning before producing the final score.
3.3.2 Stage2: Cognitive Pattern Reconstruction
Through self multi-agent rejection sampling using the Proposer–Solver–Judge pipeline, we obtain a batch of high-quality prompt–image pairs. While these pairs reflect a mapping from abstract conceptual spaces to high-dimensional visual manifolds, directly optimizing this cross-domain alignment remains stochastic and inefficient, often leading to mode collapse Chen et al. (2025a); Zhenyu et al. (2024). To move beyond this "black-box" optimization, we draw inspiration from metacognitive theory Dunlosky and Metcalfe (2008), which identifies monitoring, evaluation, and regulation as the pillars of robust learning. Based on this insight, we propose a tripartite data architecture that reclaims and structures the overlooked trajectories from the self-play cycle. By replaying these latent interactions as explicit caption, judgement, and reflection patterns, we respectively ground abstract concepts in visual features, provide evaluative signals, and encode self-correction processes. This design transforms the previously discarded internal "inner monologue" into a structured supervisory signal, fostering cognitive symmetry without external intervention.
Caption To establish robust semantic grounding, this pattern ensures the model internalizes the conceptual essence of its own creations by optimizing the inverse mapping $\pi_{\theta}(T\mid I^{*})$. By treating the highest-scoring image $I^{*}$ as the input and its originating prompt $T$ as the ground truth, the model learns to anchor abstract concepts within the specific visual manifolds it is capable of synthesizing, thereby reinforcing the bidirectional cognitive symmetry between internal concepts and external manifestations.
Judgement This pattern focuses on evaluative calibration to refine the model’s internal value system. We train the model to predict the evaluative signal $J$ for any generated pair, formulated as $\pi_{\theta}(J\mid T,I)$. By leveraging the task-specific rubrics and reasoning traces provided by the Judge, the model develops an acute perception of the latent gap between its current output and the ideal objective, providing a critical diagnostic signal for stabilizing the generative process.
Reflection Inspired by Reflexion Shinn et al. (2023), this pattern introduces iterative regulation to enhance the model’s capacity for self-evolution. Leveraging the Solver’s multiple rollouts $\{I_{1},\dots,I_{n}\}$, we utilize the rewards assigned by the Judge to identify pairs of contrasting quality, specifically selecting a high-reward "winning" image $I^{*}$ and a lower-reward "losing" image $I_{lose}$ from the same prompt. We then construct reflection trajectories formulated as $\pi_{\theta}(I^{*}\mid T,I_{lose},J)$, which explicitly encode the transition from suboptimal states to superior ones. By learning to transform the lower-quality manifestation $I_{lose}$ into its optimized counterpart $I^{*}$, the model internalizes a mechanism for self-correcting generative errors, effectively mitigating mode collapse without the need for external supervision.
These three data types are combined with high-quality self-sampled T2I generation data to fine-tune the UMM. Note that the whole reconstruction procedure is rule-based and does not introduce any complexity. Detailed generation pipeline and examples can be found in Appendix A.
3.4 UniCycle
To assess whether internal collaboration yields genuine multimodal intelligence rather than task-specific performance gains, we introduce UniCycle, a cycle-consistency benchmark that measures information preservation under a Text $\rightarrow$ Image $\rightarrow$ Text loop. Given an instruction, UniCycle evaluates whether a unified multimodal model can recover instruction-critical semantics from its own generated image through subsequent visual understanding.
  Model TIIF WISE OneIG-EN CompBench DPG Geneval
Short Long Physics Chemistry Overall Text Alignment Overall Numeracy 3d Spatial Overall Score Score
Generation Only Models
SD3 Medium 64.8 64.8 47.0 29.0 42.0 40.7 80.6 42.8 72.8 77.8 84.3 84.1 74
FLUX.1 dev 66.2 66.7 51.0 35.0 50.0 52.3 78.6 43.4 75.3 76.4 83.1 83.8 82
Unified Multimodal Models
Janus-Pro 65.4 61.1 42.0 26.0 35.0 0.1 55.3 26.7 56.4 76.2 74.0 84.3 80.0
show-o2 62.8 63.9 63.0 49.0 61.0 0.2 81.7 30.8 69.7 88.6 82.8 86.1 76.0
BLIP3-o 58.8 58.7 63.0 37.0 52.0 1.3 71.1 30.7 71.7 81.7 84.7 80.7 84.0
OmniGen2 70.2 70.3 52.0 34.0 48.0 68.0 80.4 47.5 72.0 82.2 85.8 83.6 80.0
TwiG - - - - - - - - 61.9 38.9 - - -
T2I-R1 67.6 68.3 55.0 30.0 54.0 7.3 80.4 27.7 83.3 79.4 81.9 - 77.0
BAGEL 71.0 71.8 57.0 43.0 50.0 24.4 76.9 36.1 70.4 78.0 82.2 84.0 78.0
UniCorn 74.7 72.9 67.0 47.0 55.0 46.8 84.1 42.6 83.5 84.1 88.5 86.8 82.0
$\triangle$(vs. BAGEL) +3.7 +1.1 +10.0 +4.0 +5.0 +22.4 +7.2 +6.5 +13.1 +6.1 +6.3 +2.8 +4.0
Table 1: Evaluation results on TIIF, WISE, OneIG-EN, CompBench, DPG, and Geneval benchmarks. Arrows (↑) denote that higher is better. Bold indicates the best performance across all models, and the second best is underlined. The WISE score is normalized to a 0–100 scale for visualization. Detailed comparison is listed in Appendix E.1.
Based on TIIF Wei et al. (2025), we generate QA pairs to probe instruction-implied attributes grounded in the generated image, extending the original TIIF benchmark from the T2I setting to the Text-to-Image-to-Text (T2I2T) setting. After annotation, we obtain 1,401 TIIF-style instances that cover more than ten task categories and span multiple question formats, including multiple-choice, binary (yes/no), and open-ended questions.
For evaluation, given a prompt $T$, the model first generates an image and then answers each question $q_{k}$ independently conditioned on the generated image. An external judge model assesses whether each predicted answer $\hat{y}_{k}$ is consistent with the initial prompt $T$ and the reference answer $a_{k}$, and produces a score for each question.
We define a unified metric to quantify this T2I2T consistency.
Let $\mathcal{Q}(T)$ denote the set of questions associated with a prompt $T$. We define
$\mathrm{Soft}(T)$ $=\frac{1}{|\mathcal{Q}(T)|}\sum_{k\in\mathcal{Q}(T)}s_{k},$ (2)
$\mathrm{Hard}(T)$ $=\mathbbm{1}\!\left[\forall k\in\mathcal{Q}(T),\;s_{k}=1\right].$
where $s_{k}$ denotes the judge score for question $q_{k}$, defined as a binary indicator for non-text questions and as the proportion of correctly recovered Keywords to enable a more fine-grained and continuous metric for text-type questions.
The final Soft and Hard scores are obtained by averaging over all prompts. Additional details on data construction and evaluation prompt templates are provided in Appendix D.
4 Experiments
In this section, we first introduce the experiment setup, and conduct extensive experiments to demonstrate the effectiveness of our method.
4.1 Experiment Setup
Implementation
We adopt BAGEL Deng et al. (2025b) as the base model for our main experiments. The Proposer generates 5,000 prompts, then the Solver rolls out 8 times for each prompt. Training is conducted for 600 steps on 8 NVIDIA H800 GPUs for about 7 hours with a constant learning rate of $1\times 10^{-5}$. Additional details are provided in the Appendix  A.
Figure 5: Qualitative comparison between UniCorn, BAGEL and UniCorn’s adifferent data settings. Our method jointly balances visual aesthetics, prompt fidelity, and realism in generation.
Baselines and Benchmarks
To validate our method, we compare it against three categories of approaches. First, we consider baseline models, including T2I frameworks: SD3 Medium Esser et al. (2024), FLUX.1-dev Labs (2024) and unified multimodal models: Janus-Pro Chen et al. (2025d), Show-o2 Xie et al. (2025c), BLIP3-o Chen et al. (2025b), UniGen Tian et al. (2025), TwiG Guo et al. (2025c) and T2I-R1 Jiang et al. (2025).
Figure 6: Visualization results of UniCorn at 1024×1024 resolution.
Regarding evaluation, we focus on TIIF Wei et al. (2025), WISE Niu et al. (2025), OneIG-EN Chang et al. (2025), CompBench Kil et al. (2024), DPG Hu et al. (2024), and Geneval Ghosh et al. (2023) to assess generation performance. To evaluate understanding, we further report results on standard benchmarks including MME Fu et al. (2023), MMB Liu et al. (2024b), MMMU-val Yue et al. (2024), MMVP Tong et al. (2024), and MMStar Chen et al. (2024c).
4.2 Main Results
As shown in Tab. 1, UniCorn achieves highly competitive performance across five T2I benchmarks. Our method significantly enhances fine-grained instruction following on TIIF, particularly improving robustness to short prompts (+3.7 points). On the comprehensive OneIG benchmark, UniCorn yields a 6.5-point overall improvement, with a remarkable 22.4-point gain in the Text subtask, indicating superior internalization of underlying knowledge. Furthermore, UniCorn achieves a 5 point gain on the knowledge-intensive WISE benchmark and a 6.3 point boost on CompBench. Notably, the substantial improvements in Numeracy (+13.1) and 3D Spatial (+6.1) tasks demonstrate the effective transfer of structured understanding into faithful synthesis, with UniCorn even surpassing GPT-4o on DPG benchmark (86.8 vs 86.2). These results consistently demonstrate that our self-play framework enables UMMs to bridge the gap between multimodal understanding and controllable generation, achieving robust performance that rivals state-of-the-art closed-source models.
4.3 Ablation Study
This section conducts ablation studies on data pattern, model architecture, and dataset size to further analyze our method.
4.3.1 Data Pattern
Setting TIIF-S TIIF-R MME-P MME-C MMB MMMU MMVP MMStar
Base 71.0 70.7 1685.0 696.0 84.6 52.8 69.3 65.0
Ours 74.7 78.4 1660.0 677.0 84.1 53.8 70.0 65.0
w.o. CJR 72.3 74.0 311.0 92.0 24.3 23.0 7.10 21.0
w.o. R 73.8 75.9 1632.0 655.0 84.2 53.3 71.3 65.0
w.o. J 74.2 74.8 1542.0 478.0 82.6 51.9 65.3 61.0
w.o. C 74.5 76.4 1653.0 701.0 84.3 50.9 68.0 64.0
w.o. G 73.4 72.3 1669.0 685.0 84.2 53.0 70.0 64.0
Table 2: Ablation study on data composition. Each variant is trained independently by removing exactly one data type from the full GCJR setting while keeping all other components fixed. S and R denote Short Score and Real Score, respectively.
Model Hard score
Bagel 36.6
Show-o2 36.1
Janus-Pro 9.9
UniCorn* 40.0
UniCorn 46.5
Table 3: Hard score results on UniCycle. Soft score results are reported in Appendix 17.
This section deconstructs multimodal data patterns to demonstrate how Cognitive Pattern Reconstruction bridges the gap between understanding and generation within a unified framework.
Tab. 2 reveals a hierarchical synergy between data patterns: while relying solely on generation (w.o. CJR) maintains basic instruction following (TIIF-S: 72.3), it triggers a catastrophic collapse of the latent space, evidenced by the sharp drop in MME-P from 1685.0 to 311.0. This proves that unconstrained generative training without semantic grounding leads to mode collapse. Conversely, incorporating Cognitive Pattern Reconstruction patterns (C, J, R) stabilizes the model; Judgment and Reflection provide evaluative signals that boost complex generative quality (TIIF-R: 78.4), while Captioning preserves the multimodal foundation and spatial reasoning capabilities. Finally, although removing generation (w.o. G) maintains comprehension metrics like MME-P (1669.0), it stalls generative growth, resulting in lower TIIF scores (73.4). Qualitative comparisons are shown in Fig. 5. Ultimately, these results confirm a reciprocal reinforcement: generative trajectories reconstructed as interpretive signals refine semantic boundaries, which in turn guide higher-fidelity synthesis, allowing UniCorn to significantly improve generation while preserving its core multimodal intelligence.
4.3.2 Model Architecture
We first demonstrate the effectiveness of our method on BAGEL, where understanding and generation components are decoupled. To evaluate its generalization to tightly coupled architectures, we conduct a base model ablation on the purely autoregressive Janus-Pro-7b Chen et al. (2025d). Tab. 4 shows that our method improves Janus by +3.2 on TIIF, +7.0 on WISE, and +4.7 on OneIG-EN, with the most pronounced gain on the knowledge-intensive WISE benchmark. This suggests that the proposed approach enhances knowledge expression by guiding generation through improved understanding, a mechanism that generalizes across different model architectures.
4.3.3 Scaling Law for UniCorn
Scaling laws guide architectural design and optimization Kaplan et al. (2020); Chen et al. (2024e), but prior methods scale poorly due to reliance on external models or fixed prompts. In contrast, UniCorn achieves scalable self-improvement through unbounded self-sampling and efficient Cognitive Pattern Reconstruction. To explicitly assess scalability, we conduct scale-up experiments by varying the amount of self-generated data across {1k, 5k, 8k, 10k, 20k}.
Model TIIF+0.0 WISE(+0.0) OneIG-EN+0.0
Janus Pro 63.2+0.0 35.0+0.0 26.7+0.0
+UniCorn 65.9+2.7 42.0+7.0 31.4+4.7
UniCorn 73.8+0.0 55.0+0.0 42.6+0.0
UniCorn* 74.4+0.6 54.0-1.0 44.9+2.3
Table 4: Ablation studies on the base model (top) and the self-play framework (bottom). Unicorn denotes the BAGEL model trained on data constructed using Qwen3-VL-235B-A22B-Instruct.
As shown in Fig. 7, with only 1K training samples, our method already surpasses RecA  (Xie et al., 2025b) on TIIF. As the data scale increases, the model’s generative performance continues to improve, with more pronounced gains on long-prompt generation; notably, with just 5k samples, it outperforms IRG Huang (2025) trained on 30k GPT-4o distilled data as well as the strong closed-source model DALL·E 3 Betker et al. (2023a). These results reveal a favorable scaling regime in which self-generated data alone suffices to drive continual and efficient improvements in generative capability.
4.4 Analysis
We design a series of experiments for in-depth analysis to address the following two questions.
Q1: Is self-play necessary?
Figure 7: Data scaling result on TIIF. The score consistently improves when the dataset size scales up. Notably, UniCorn surpasses many powerful models only using 5k training data.
For self-play assessment, we use Qwen3-VL-235B-A22B-Instruct Yang et al. (2025) for data construction (UniCorn*). As shown in Tab. 4, employing stronger proposers/judges yields diminishing returns, where high costs and training time outweigh performance gains. This likely stems from the difficulty of fitting high-entropy teacher distributions, increasing latency without proportional information gain. We then compare UniCorn with four unified models on UniCycle (Judge: Qwen3-235B-A22B Yang et al. (2025)). UniCycle requires both generation and understanding, reducing task bias and evaluating the model’s self-reflection, thus signaling comprehensive multimodal intelligence.
As shown in Tab. 3, UniCorn achieves the highest Hard score (46.5), outperforming its base BAGEL by nearly 10 points and others by over 3 points. UniCorn* lags by 6.5 points, suggesting strong external supervision yields disproportionate costs and insufficient unified coordination. This demonstrates that self-play enhances unified capabilities by distilling understanding into generation without degradation, achieving SOTA on UniCycle. In contrast, Janus-Pro significantly underperforms comparable-scale models on UniCycle, revealing a gap between its generation and self-understanding.
Q2: Why UniCorn works?
UniCorn addresses the asymmetry in Unified Multimodal Models where strong understanding capabilities remain inactive during generation. We identify three critical limitations: (1) the model lacks a holistic perception of the content it is about to generate, (2) it does not actively assess the quality of its own outputs during generation, and (3) it lacks the ability to reflect on and revise suboptimal generations. UniCorn resolves this by enabling understanding to supervise generation through a unified self-improvement loop involving captioning, evaluation, and reflection, restoring alignment for more faithful and controllable results.
Theoretically, we justify this approach using Mutual Information and Bayes’ theorem, demonstrating that our task decomposition effectively minimizes Negative Log Likelihood (NLL). This guarantees that the auxiliary understanding tasks mathematically optimize the final unified objective. Detailed derivation is presented in Appendix C.
5 Conclusion
In this paper, we propose UniCorn, a self-supervised post-training framework that unifies multimodal comprehension and generation within a single model via multi-agent self-play and Cognitive Pattern Reconstruction, distilling internal latent knowledge into high-quality generative signals without external supervision. Extensive experiments, including the UniCycle cycle-consistency benchmark, demonstrate significant improvements in T2I generation while preserving multimodal intelligence, highlighting self-contained feedback loops as a scalable path for unified multimodal models.
Limitations
Despite achieving robust performance in both T2I generation and multimodal understanding, UniCorn possesses certain limitations. First, the current self-improvement framework operates in a single-turn manner and primarily enhances generative capabilities, with no significant gains observed in understanding metrics. In future work, we intend to explore multi-turn iterative self-play to foster the co-evolution of both capabilities. Second, the self-play mechanism requires the UMM to handle prompt generation, rollout, and judgment, which inevitably introduces additional computational costs. We plan to investigate more efficient methodologies to streamline this process in subsequent research.
Ethical Statement
The development of UniCorn adheres to ethical standards for AI research. We utilize publicly available open-source models as our foundation and conduct all experiments using standard public benchmarks. Our self-improvement framework aims to enhance generative quality through internal feedback, thereby reducing the need for massive external data collection. While we implement internal filters during the self-play process to improve output alignment, we acknowledge that multimodal models may still reflect biases present in their pre-training data. We are committed to transparency and encourage the responsible use of our framework in downstream applications.
Appendix A Experiment Details
Data
The specific prompt category, judgement rubrics for § 3.3.1 is shown in Tab 7. We set different random seeds and cfg_text_scale in image sampling for a single prompt, to increase diversity for better images. Moreover, to ensure data quality, we filter the groups of samples when the highest score produced by the Judge is less than a fixed threshold (we choose 7).
For § 3.3.2, we use dozens of hand-written templates, without increasing any computational complexity. For Caption data, note that some image generation prompts contain phrases like "generate an image" or "create an image. For these types of data, we transfer the traditional caption task into "reconstruct image generation prompt, serving as a generalized caption task, which enhances data diversity and maintains data quality.
The data mixture we use is 5k for Generation, 5k for Caption, 3k for Judgement, and 1k for Reflection. The detailed examples for training data are shown in Tab. 8.
Training
We conduct the post-training phase using the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-15}$) with a constant learning rate of $1\times 10^{-5}$. To ensure training stability, we implement a warm-up of 50 steps and apply gradient clipping at a threshold of 1.0. We set the total training duration to 600 steps, utilizing a gradient accumulation of 4 steps to manage the effective batch size. Furthermore, we apply an EMA ratio of 0.99 and balance the training objective with a CE to MSE loss weight ratio of $0.1:1$. For task-specific configurations, we utilize a maximum context window of 40k tokens, with generation and understanding resolutions set to $(512,1024)$ and $(378,980)$, respectively. Finally, a diffusion timestep shift of 4.0 is applied to calibrate the generative process. We conduct all experiments on 8 NVIDIA H800 (80 GB) GPUs. Tab. 5 shows the detailed hyperparameter configurations when post-training BAGEL.
Benchmarks
To evaluate the generative performance of our model, we employ six representative text-to-image (T2I) benchmarks that assess various dimensions of synthesis quality and semantic alignment:
  • TIIF: This benchmark evaluates the model’s ability to follow complex prompts across different lengths, specifically categorized into TIIF-S (short) and TIIF-L (long) to measure fine-grained text-to-image alignment. We use the official testmini subset and choose QwenVL2.5-72B Bai et al. (2025) as the evaluators.
  • WISE: This metric focuses on spatial consistency and visual fidelity, utilizing normalized scores to reflect the model’s performance in complex scene layout generation.
  • OneIG: A large-scale generative benchmark designed to test the robustness and diversity of the model across a wide array of semantic categories.
  • CompBench: This benchmark targets compositional generation, specifically assessing how well the model handles attribute binding, object relations, and numerical constraints.
  • DPG : DPG emphasizes the reconstruction of dense, multi-entity prompts, requiring the model to accurately synthesize multiple subjects and their respective fine-grained attributes.
  • GenEval: A comprehensive evaluation framework that employs automated metrics to quantify core generative capabilities, including object recognition and attribute alignment.
Together, these benchmarks provide a multi-dimensional assessment of our model’s capacity to transform abstract conceptual prompts into high-fidelity visual outputs while maintaining strict adherence to textual constraints.
Hyperparameters Post-training
Learning rate $1\times 10^{-5}$
LR scheduler Constant
Weight decay 0.0
Gradient norm clip 1.0
Gradient accumulation steps 4
EMA ratio 0.99
Loss weight (CE: MSE) 0.1: 1
Optimizer AdamW ($\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-15}$)
Warm-up steps 50
Training steps 600
Max context window 40k
Gen resolution (min short side, max long side) (512, 1024)
Und resolution (min short side, max long side) (378, 980)
Diffusion timestep shift 4.0
Table 5: Training recipe of UniCorn.
Appendix B More Related Work
Unified Multimodal Models
UMMs aim to unify cross-modal understanding and generation, yet a persistent challenge is that strong understanding does not reliably translate into equally strong native generation. Most UMMs follow two main architectural routes. Pure autoregressive models jointly predict text and visual tokens with a unified next-token objective over interleaved sequences, as in Janus-Pro, Emu, and MetaMorph Chen et al. (2025e); Cui et al. (2025); Tong et al. (2025). Hybrid designs keep autoregressive modeling for language while relying on diffusion-style synthesis for continuous images, either via integrated modeling within a single backbone (e.g., Show-o and MonoFormer Xie et al. (2024); Zhao et al. (2024)) or through modular routing and sparse experts (e.g., LMFusion, Mixture-of-Transformers, and BAGEL Shi et al. (2024); Liang et al. (2024b); Deng et al. (2025a)), with related paradigms such as Diffusion Forcing exploring diffusion-style guidance for interleaved generation Chen et al. (2024a); Huang et al. (2025c). Recent work also explores image generation foundation models that systematize text-conditioning and training recipes on diffusion-style backbones. Qwen-Image Wu et al. (2025a) adopts a double-stream MMDiT, conditioned on a frozen Qwen2.5-VL encoder and a VAE tokenizer, and uses progressive/curriculum training with multi-task objectives to cover settings such as multilingual text rendering and editing. Z-Image Cai et al. (2025) proposes a 6B single-stream diffusion Transformer (S3-DiT) and derives a Turbo variant via few-step distillation and reward post-training, focusing on inference efficiency under low sampling steps and text-rendering-related scenarios. Beyond architecture, recent work investigates self-improvement by turning self-produced signals into training objectives Yu et al. (2025); Zhou et al. (2024b); Wang et al. (2025b). For UMMs, SRUM leverages internal understanding as an evaluator to derive fine-grained rewards Jin et al. (2025a), and UniRL couples generation and understanding via supervised and reinforcement learning Mao et al. (2025). Complementary directions also study data-centric enhancement for vision-language alignment, e.g., ultra-detailed caption generation to enrich training signals for VLMs Zeng et al. (2025b), and interaction-based learning setups that emphasize perception and reasoning behaviors Zeng et al. (2025a). However, existing self-improvement pipelines often depend on auxiliary components or externally produced dense feedback, as well as task-specific reward shaping, fixed prompt pools, or pre-selected concepts, which can limit scalability and generalization when extending self-improvement to broader UMM settings.
Self-Improvement for LLMs
Self-play drives autonomous improvement by pairing self-generated challenges with outcome-driven learning Silver et al. (2017). In LLMs, this enables zero-data self-evolution: models generate training signals without curated datasets, as exemplified by the uncertainty and self-consistency curricula of R-Zero Huang et al. (2025a) as well as the executor-verified rewards of Absolute Zero Zhao et al. (2025a). Beyond task generation, self-produced evaluation guides preference learning and reasoning, through methods such as self-rewarding feedback Yuan et al. (2024), constraint-based optimization Zhou et al. (2024b), reflective reward learning Choi et al. (2024), and process-consistency rewards for long-horizon tasks Guo et al. (2025b). In multimodal settings, related efforts incorporate retrieval-augmented reasoning with reinforcement learning to improve understanding over visually rich information sources Wang et al. (2025a). Extending to Unified Multimodal Models (UMMs), their integrated modules naturally enable self-improvement, with the understanding module providing internal multi-scale feedback to guide generation, establishing a promising paradigm for fully model-driven enhancement.
Multi-Agent Systems for LLMs
LLM-based multi-agent systems instantiate role-specialized agents to decompose tasks, explore diverse solutions, and cross-check results, supported by orchestration frameworks such as AGENTVERSE Chen et al. (2024d) and debate-style interactions that encourage diverse hypotheses and mutual critique Liang et al. (2024a). While many systems are primarily deployed at inference time, recent work explores closed-loop training with self-generated signals, including multi-agent training and self-play pipelines Motwani et al. (2024); Zhao et al. (2025a); Chen et al. (2025f). Beyond pure language agents, embodied and action-centric extensions seek to improve generalization by structuring the coupling between reasoning and action, as in DualVLA Fang et al. (2025). However, empirical analyses show these systems can be brittle and costly, with recurring coordination and verification failures that limit scalability and generalization Cemri et al. (2025). Motivated by these limitations, our work uses lightweight role instantiation within a single unified multimodal model, turns role interactions into self-training signals for improving native multimodal generation, and introduces a cycle-consistency benchmark to test whether gains reflect genuine multimodal understanding rather than task-specific tuning.
LLM-as-a-Judge
Recent work increasingly adopts LLM-as-a-Judge as a scalable alternative to human evaluation for open-ended generation and benchmark construction, where strong LLMs provide pointwise scores or comparative rankings with broader coverage than heuristic metrics at lower annotation cost Li et al. (2025, 2024b). However, LLM-based judges are not uniformly reliable. Their judgements can be sensitive to prompt phrasing and candidate presentation, and they may exhibit systematic biases and vulnerabilities, including adversarial manipulation Raina et al. (2024); Thakur et al. (2025); Li et al. (2024b). These concerns motivate meta-evaluation of judges and evaluation protocols that reduce reliance on fragile or implicit judgement signals Feng et al. (2025), as well as targeted benchmarks that stress-test self-critique in tool-calling error scenarios Huang et al. (2025b). In multimodal evaluation, video reasoning benchmarks further broaden coverage beyond static images, including chain-of-thought video reasoning and visual-prompt-based interaction protocols Qi et al. (2025); Zhao et al. (2025c). In our setting, judge models serve two roles. For T2I generation, we use a VLM-based judge with task-specific rubrics to assess prompt-image alignment and visual fidelity. For T2I2T evaluation, we use the same UMM as an LLM-based judge to verify whether predicted answers match the original instruction and reference answers, enabling structured scoring at scale.
Appendix C Theoretical Analysis
In this section, we present a thorough theoretical analysis to explain why UniCorn works. As discussed in § 3.3.2 and Appendix A, we construct the following four types of data:
  • Generation Data ($G$): High-quality images sampled by the model and selected via a Best-of-N strategy.
  • Captioning Data ($C$): Constructed via a reverse process, where the best images and caption prompts serve as input to predict the original generation prompts.
  • Judgement Data ($J$): Self-evaluation outputs from the model, including Chain-of-Thought (CoT) reasoning and final scoring.
  • Reflection Data ($R$): Self-correction data taking suboptimal images and editing instructions as input to output the optimal images.
Ablation studies demonstrate that each data type contributes to both generation and understanding capabilities, fostering a truly unified model. Below, we provide a theoretical analysis of why these four synthetic data types synergistically enhance image generation.
For most unified multimodal model like BAGEL, parameters for generation and understanding are shared partially. The objective is to learn the joint distribution of Text $(T)$ and Image $(I)$, denoted as $\pi_{\star}(T,I)$. We train the model $\pi_{\theta}(T,I)$ to approximate the constructed data distribution $p(T,I),(T,I)\sim\mathcal{D}$, where $\mathcal{D}$ is the predefined dataset, by minimizing the unified loss function $\mathcal{L}_{Unified}$.
C.1 Bi-Directional Mutual Information
Most existing works focus solely on constructing $p(I\mid T)$ to enhance generation. However, our experiments show that this single-directional training leads to a collapse in understanding capabilities. We analyze this from the perspective of Mutual Information.
Consider the mutual information between image and text, $MI(I;T)$
$\begin{split}MI(I;T)&=H(I)-H(I\mid T)\\ &=H(T)-H(T\mid I)\end{split}$ (3)
Constructing only generation data $p(I\mid T)$ minimizes an upper bound on $H(I\mid T)$(the conditional cross-entropy/NLL). However, according to the equation above, one-way likelihood training provides no direct training signal for the other. The model fails to capture the dependency of $T$ given $I$, leading to the collapse of understanding capabilities. Due to parameter sharing, this representational deficiency results in sub-optimal generation performance.
By constructing Captioning Data $p(T\mid I)$ via a self-dual approach, we encourage bidirectional consistency between the two conditionals:
$\begin{split}p(I,T)&=p(I\mid T)\,p(T)\\ &=p(T\mid I)\,p(I),\quad(T,I)\sim\mathcal{D}_{C}\end{split}$ (4)
where $\mathcal{D}_{C}$ is the caption dataset we constructed, and $p(I)$, $p(T)$ are priors determined by both the dataset and model.
This explains why Caption data not only preserves understanding capabilities but also enhances generation by enforcing a more robust, bi-directionally consistent multimodal representation.
C.2 Internalized Preference Judgement
A truly unified multimodal model requires not only the ability to generate and understand but also the capacity to align with human preferences—specifically, the ability to Judge. We refine the target distribution to include judgement $J$ J, denoted as $\pi_{\star}(T,I,J)$. Using the chain rule of probability, the model’s joint distribution can be decomposed as
$\pi_{\theta}(I,T,J)=\pi_{\theta}(J\mid I,T)\cdot\pi_{\theta}(I\mid T)\cdot\pi_{\theta}(T)$ (5)
This decomposition implies that the system is composed of text priors, text-to-image generation, and the ability to judge the quality of the $(I,T)$ pair. We construct judgement Data $p(J\mid I,T)$ to train the term $\pi_{\theta}(J\mid I,T)$. This allows the model to "internalize" evaluation metrics, effectively learning a discriminator that guides the generator toward higher-quality outputs.
C.3 Trajectory of Self-Reflection
With the introduction of judgement, the model can learn to improve from "bad" to "good" states. We aim for the model to generate the optimal image $I_{w}$ potentially via an iterative process.
Let $I$ denote a suboptimal image sampled during exploration. We can model the generation of the best image $I^{*}$ by introducing $I$ as an intermediate latent variable in the probability decomposition:
$\pi_{\theta}(I^{*}\mid T,J)=\pi_{\theta}(I^{*}\mid I,T,J)\cdot\pi_{\theta}(I\mid T,J)$ (6)
Here, $\pi_{\theta}(I\mid T,J)$ represents the initial generation, and $\pi_{\theta}(I^{*}\mid I,T,J)$ represents the refinement step.By constructing Reflection Data $p(I^{*}\mid I,T,J)$, we explicitly train the model to act as a "correction operator". This enables the model to learn the trajectory of improvement, significantly boosting its ability to handle complex instructions and self-correction.
C.4 Objective Decomposition for Unified Multimodal Learning
From the perspective of Negative Log-Likelihood (NLL), the overall loss function $\mathcal{L}_{Unified}$ for BAGEL is decomposed as follows:
$\mathcal{L}_{Unified}=\mathcal{L}_{G}+\mathcal{L}_{C}+\mathcal{L}_{J}+\mathcal{L}_{R},$ (7)
where $\lambda_{i}$ represents the relative data proportions across each dataset. The individual loss components are defined as:
$\mathcal{L}_{G}$ $=-\mathbb{E}_{(I^{*},T)\sim\mathcal{D}_{bon}}\left[\log\pi_{\theta}(I^{*}\mid T)\right]$
$\mathcal{L}_{C}$ $=-\mathbb{E}_{(T,I^{*})\sim\mathcal{D}_{C}}\left[\log\pi_{\theta}(T\mid I^{*})\right]$
$\mathcal{L}_{J}$ $=-\mathbb{E}_{(I,T,J)\sim\mathcal{D}_{J}}\left[\log\pi_{\theta}(J\mid I,T)\right]$
$\mathcal{L}_{R}$ $=-\mathbb{E}_{(I^{*},I,T,J)\sim\mathcal{D}_{R}}\left[\log\pi_{\theta}(I^{*}\mid I,T,J)\right]$
where $\mathcal{D}_{i}$ represents different synthetic datasets.
Appendix D Benchmark Details
D.1 Data Construction
Based on the TIIF benchmark, we generate question–answer pairs for instruction reconstruction, extending the original Text-to-Image (T2I) evaluation to a Text-to-Image-to-Text (T2I2T) setting. To balance task difficulty with answer evaluability, we design task- and question-type–specific prompt templates. For negation-related tasks, we construct prompts that elicit binary (yes/no) questions, ensuring unambiguous evaluation. For tasks like spatial relation, open-ended questions often lead to ambiguous judgments— for instance, an instruction specifies "left" but the generated image places an object in the "front-left" position, an answer such as "in front" may be plausible yet difficult to assess consistently. To improve evaluation stability, we therefore formulate these tasks as multiple-choice questions. For the remaining task types, such as color recognition and counting, we adopt open-ended question–answer formats to maintain sufficient difficulty and discriminative power. Moreover, we explicitly enforce task-type–based question completeness: since all instruction-implied information relevant to the task type is treated as a reconstruction target, a QA set is considered valid only if it fully covers the reconstruction targets without redundancy. QA pairs are generated using Qwen3-235B-A22B Yang et al. (2025) and subsequently annotated under a unified labeling protocol by experienced human annotators. After filtering, we retain 1,401 high-quality instances and totally 2968 questions ( The distribution of question types is shown in Tab.16) for evaluation, covering almost all task types present in the original TIIF benchmark. We present several cases in Fig. 10.
D.2 UniCycle Evaluation Prompt
The prompt templates for T2I2T evaluation of UniCycle are presented in Fig. 11, 12.
D.3 Soft scores results on UniCycle
Soft scores results of UniCorn and the other four models on UniCycle are shown in Tab. 17.
Appendix E Additional Results
E.1 Quantitative Experiments
Detailed scores across the six T2I benchmarks are reported in Tab. 9,  10,  111213and  14. We also evaluate our model on the image edit task Zhao et al. (2025b) in Tab. 6.
Model RISE Score
BAGEL 33.33
UniCorn 38.87(+5.54)
Table 6: The evaluation results of RISE.
E.2 Qualitative Results
The qualitative comparison of the reliance on external data and models between our approach and other methods is presented in Tab. 15. Without relying on external task-specific models or annotated data, UniCorn achieves state-of-the-art performance on OneIG-EN using only 5K training samples.
E.3 Failure Cases
In Fig. 13, we show two failure cases of UniCorn in challenging tasks such as Negation and Counting. We attribute the model’s limitations on these tasks to their inherent difficulty for multimodal models. Within our self-play training paradigm, it is challenging to provide effective supervision for such tasks; consequently, the lack of significant improvement is consistent with our expectations.
Major Category Generation Requirement Judgement Rubrics Example
General Object Depict specific real-world objects or scenes, focusing on attributes including shape, color, texture, and single/multi-object composition. Object existence, attribute accuracy (color/shape/texture), and compositional correctness. [Uncaptioned image]
Object Relations Reflect logical connections between objects, involving action & interaction, comparison, differentiation, or negation. Logical correctness of relations (e.g., A is interacting with B), in addition to basic object correctness. [Uncaptioned image]
General Knowledge Depict specific general elements requiring external knowledge in real life, such as festivals, sports, celebrities, religions, or crafts. Factual accuracy, cultural recognition, and attribute alignment with real-world entities. [Uncaptioned image]
Spatio Reasoning Handle complex spatial layouts, including 2D/3D structures, occlusion reasoning, and specific viewing perspectives (e.g., bird’s-eye view). Spatial consistency, perspective correctness, and accurate handling of occlusions/depth. [Uncaptioned image]
Temporal Reasoning Reflect time-sensitive states, such as horizontal time (synchronous elements) or longitudinal time (chronological changes/stages). Temporal consistency, logical progression of states, and accuracy of time-specific features. [Uncaptioned image]
Text Rendering Render legible text within images across various formats: natural-scene text, designed posters/menus, or handwriting/graffiti. OCR accuracy (spelling), font style appropriateness, and text-background integration. [Uncaptioned image]
Natural Science Accurately represent scientific subjects, including precise anatomy of animals/plants and physics/chemistry phenomena. Scientific realism, biological anatomical correctness, and physical plausibility. [Uncaptioned image]
Portrait Generate human-centered portraits with specific framing requirements: close-up, half-body, or full-body shots. Framing accuracy (shot scale), facial/anatomical correctness, and identity consistency. [Uncaptioned image]
Stylization Adhere to specific artistic styles, primarily focusing on Anime style or various artistic stylizations (e.g., oil painting, sketch). Style fidelity, aesthetic quality, and texture application consistent with the requested art medium. [Uncaptioned image]
Counting Generate a precise number of specific objects as described in the prompt. Count accuracy (numerical precision) and object distinctness. [Uncaptioned image]
Table 7: Detailed data type range, description and judgement rubrics.
Category Prompt Example Response
Generation A glass sculpture in the shape of a turtle with intricate patterns of red lines on its shell, resting on a black marble pedestal, with soft light coming from above, highlighting the contours of the turtle and casting delicate shadows on the floor. [Uncaptioned image]
Caption Type1: Infer the image-generation prompt from the picture.
Type2:Please provide a caption for this picture.
A glass sculpture in the shape of a turtle with intricate patterns of red lines on its shell, resting on a black marble pedestal, with soft light coming from above, highlighting the contours of the turtle and casting delicate shadows on the floor.
Judgement Judge the image quality based on the generation prompt. (See Fig. 9)
{ "analysis": "The generated image accurately depicts a glass sculpture in the shape of a turtle with intricate patterns of red and blue lines on its shell……(Lengthy analysis text omitted)", "score": 10}
Reflection A glass sculpture in the shape of a turtle with intricate patterns of red lines on its shell, resting on a black marble pedestal, with soft light coming from above, highlighting the contours of the turtle and casting delicate shadows on the floor. [Uncaptioned image] $\longrightarrow$ [Uncaptioned image]
Table 8: Examples of Generation, Caption, Judgement, Reflection training data. We choose the same image and prompt for better illustration.
Table 9: Quantitative evaluation results on OneIG-EN.
[…]
Janus-Pro (Chen et al., 2025d) 0.553 0.001 0.139 0.276 0.365 0.267
T2I-R1 (Jiang et al., 2025) 0.804 0.073 0.167 0.290 0.277 0.322
BLIP3-o (Chen et al., 2025c) 0.711 0.013 0.223 0.361 0.229 0.307
BAGEL (Deng et al., 2025a) 0.769 0.244 0.173 0.367 0.251 0.361
Show-o2-7B (Xie et al., 2025c) 0.817 0.002 0.226 0.317 0.177 0.308
SDv1.5 (Rombach et al., 2022) 0.565 0.010 0.207 0.383 0.429 0.319
SDXL (Podell et al., 2024) 0.688 0.029 0.237 0.332 0.296 0.316
FLUX.1-dev (Labs, 2024) 0.786 0.523 0.253 0.368 0.238 0.434
SD3 (Esser et al., 2024) 0.805 0.407 0.293 0.386 0.244 0.427
FLUX.1-dev (Labs, 2024) 0.786 0.523 0.253 0.368 0.238 0.434
SANA-1.5 4.8B (PAG) (Xie et al., 2025a) 0.765 0.069 0.217 0.401 0.216 0.334
Lumina-Image 2.0 (Qin et al., 2025) 0.819 0.106 0.270 0.354 0.216 0.353
IRG*  (Huang et al., 2025c) 0.839 0.377 0.239 0.427 0.192 0.415
OmniGen2  (Xiao et al., 2025) 0.804 0.680 0.271 0.377 0.242 0.475
UniCorn 0.841 0.468 0.232 0.395 0.203 0.426
GPT-4o (OpenAI, 2025) 0.851 0.857 0.345 0.462 0.151 0.533
Table 10: Quantitative evaluation results of instruct-following capability on TIIF testmini (QwenVL2.5-72B as the evaluation). * indicates that the model has not yet been open-sourced; we report the metrics as presented in the official paper.
Model Overall Basic Following Advanced Following Designer
Avg Attribute Relation Reasoning Avg Attribute +Relation Attribute +Reasoning Relation +Reasoning Style Text Real World
short long short long short long short long short long short long short long short long short long short long short long short long
FLUX.1-dev (Labs, 2024) 66.24 66.72 74.41 76.67 72.50 75.50 78.20 79.78 72.52 74.73 60.72 60.95 66.76 65.50 61.76 60.74 56.60 57.49 63.33 60.00 44.49 54.75 74.63 72.01
FLUX.1-Pro (Labs, 2024) 63.75 63.53 71.39 73.57 70.00 68.50 68.51 79.97 75.66 72.23 64.63 61.42 70.69 72.99 62.34 57.27 64.65 57.11 63.00 63.00 34.39 36.65 69.94 66.78
DALL-E 3 (Betker et al., 2023b) 74.47 72.94 77.35 78.40 77.62 75.00 80.22 79.67 74.22 80.54 70.11 68.45 76.65 75.05 68.39 68.07 63.64 59.92 79.31 80.00 74.07 75.51 76.12 62.69
SD3.5-large (Esser et al., 2024) 68.69 64.92 73.72 72.10 77.50 66.50 74.79 77.16 68.87 72.64 65.59 63.41 70.85 68.22 65.03 62.93 61.03 61.66 56.67 60.00 73.30 46.15 70.15 69.03
PixArt-$\Sigma$ (Chen et al., 2024b) 57.46 57.04 67.74 68.19 65.50 69.50 74.33 72.11 63.40 62.96 56.71 54.52 62.47 59.67 57.51 55.08 54.84 52.64 76.67 73.33 2.71 4.98 63.06 63.06
Show-o (Xie et al., 2024) 57.34 61.33 69.99 75.30 66.50 80.00 76.47 71.88 67.00 74.04 58.25 58.19 67.21 64.33 54.26 58.86 61.38 56.19 46.67 66.67 4.98 11.31 71.64 68.66
Janus-Pro-7B (Chen et al., 2025d) 65.38 61.10 74.99 73.19 74.50 78.00 73.69 70.51 76.77 71.04 61.77 56.03 65.71 66.48 62.01 55.62 61.16 49.34 43.33 70.00 38.46 42.08 79.48 73.51
T2I-R1 (Jiang et al., 2025) 67.61 68.34 81.14 79.45 80.50 78.50 83.09 79.49 79.81 80.37 67.38 65.90 69.92 65.27 70.10 71.62 68.69 64.68 50.00 63.33 32.13 37.56 74.25 74.25
BAGEL (Deng et al., 2025a) 70.97 71.79 78.16 78.12 78.00 79.50 80.24 79.08 76.25 75.77 68.23 68.19 73.37 77.49 64.36 66.15 68.92 61.48 80.00 80.00 40.72 52.40 76.87 74.63
MidJourney v7 (Midjourney, 2025) 65.92 62.43 73.96 74.63 75.00 82.00 78.74 78.51 68.12 68.55 63.44 62.59 70.60 74.03 64.43 59.58 58.84 61.34 66.67 33.33 31.67 34.39 79.22 75.32
Show-o2  (Xie et al., 2025c) 62.80 63.87 75.30 74.45 73.00 71.00 77.22 74.09 75.69 78.25 61.38 66.12 63.47 67.44 62.63 70.31 64.15 60.00 60.00 33.33 14.03 10.86 75.00 74.63
BAGEL (Deng et al., 2025a) 68.06 68.78 77.63 79.40 75.00 77.00 78.55 82.37 79.33 78.81 71.24 68.20 77.65 75.37 69.77 65.87 72.93 67.91 69.93 63.33 26.24 26.70 69.78 71.64
IRG*  (Huang et al., 2025c) 76.00 73.77 83.17 81.28 81.00 76.00 82.96 81.86 85.54 85.98 75.25 74.66 75.82 77.25 78.16 77.76 73.84 72.93 90.00 70.00 43.89 47.51 72.76 74.63
UniCorn 74.70 72.94 79.43 78.53 81.50 79.50 83.14 79.84 73.64 76.25 73.39 71.81 76.84 74.59 72.34 71.33 72.81 71.66 73.33 76.67 58.85 49.77 79.85 76.87
GPT-4o (OpenAI, 2025) 84.19 84.61 85.30 86.55 81.00 82.12 86.16 84.12 88.74 94.50 81.24 79.75 81.95 81.55 80.03 79.85 80.88 75.68 76.67 86.67 92.76 90.05 89.55 88.06
Table 11: Comparison of world knowledge reasoning on WISE. WISE examines the complex semantic understanding and world knowledge for T2I generation. ‘Gen. Only’ stands for an image generation model, and ‘Unified’ denotes a model that has both understanding and generation capabilities. * indicates that the model has not yet been open-sourced; we report the metrics as presented in the official paper.
[…]
Gen. Only SDv1.5 (Rombach et al., 2022) 0.34 0.35 0.32 0.28 0.29 0.21 0.32
SDXL (Podell et al., 2024) 0.43 0.48 0.47 0.44 0.45 0.27 0.43
SD3.5-large (Esser et al., 2024) 0.44 0.50 0.58 0.44 0.52 0.31 0.46
PixArt-Alpha (Chen et al., 2024b) 0.45 0.50 0.48 0.49 0.56 0.34 0.47
playground-v2.5 (Li et al., 2024a) 0.49 0.58 0.55 0.43 0.48 0.33 0.49
FLUX.1-dev (Labs, 2024) 0.48 0.58 0.62 0.42 0.51 0.35 0.50
Unified Janus (Wu et al., 2025b) 0.16 0.26 0.35 0.28 0.30 0.14 0.23
Show-o-512 (Xie et al., 2024) 0.28 0.40 0.48 0.30 0.46 0.30 0.35
Janus-Pro-7B (Chen et al., 2025d) 0.30 0.37 0.49 0.36 0.42 0.26 0.35
Emu3 (Wang et al., 2024b) 0.34 0.45 0.48 0.41 0.45 0.27 0.39
MetaQuery-XL (Pan et al., 2025) 0.56 0.55 0.62 0.49 0.63 0.41 0.55
BAGEL (Deng et al., 2025a) 0.42 0.53 0.64 0.42 0.57 0.43 0.50
Show-o2 (Xie et al., 2025c) 0.64 0.58 0.61 0.58 0.63 0.49 0.61
T2I-R1  (Jiang et al., 2025) 0.56 0.55 0.63 0.54 0.55 0.30 0.54
BLIP3-o  (Chen et al., 2025b) 0.49 0.51 0.63 0.54 0.63 0.37 0.52
UniCorn 0.48 0.56 0.67 0.47 0.67 0.47 0.55
GPT-4o (OpenAI, 2025) 0.81 0.71 0.89 0.83 0.79 0.74 0.80
Table 12: Comprehensive T2I-CompBench Results. This table includes T2I (Labs, 2024; Esser et al., 2024; Podell et al., 2024) and UMMs (Chen et al., 2025d; Xie et al., 2025c).
Model 3d Spatial Color Complex Nonspatial Numeracy Shape Spatial Texture Overall
T2I Models
FLUX.1-dev 76.39 90.63 83.51 87.47 75.30 80.20 84.23 87.07 83.10
FLUX.1-schnell 79.38 84.53 81.96 85.55 72.82 82.20 85.49 86.38 82.29
SD-3-medium 77.83 91.63 84.73 86.12 72.80 83.72 88.20 89.03 84.26
SD-xl-base-1 72.25 77.75 75.00 85.28 57.14 72.18 77.08 78.38 74.38
Unified Multimodal Models
Janus-Pro 76.17 84.25 80.28 80.47 56.43 65.14 79.67 69.67 74.01
T2I-R1 79.35 92.11 85.48 83.32 69.47 74.08 86.44 84.85 81.89
Show-O2 88.61 87.73 87.88 85.91 69.74 73.99 86.60 82.17 82.83
OmniGen2 82.21 92.22 86.87 88.51 72.00 83.95 90.07 90.88 85.84
BLIP3o 81.73 89.92 85.55 84.78 71.67 83.75 92.47 87.45 84.66
BAGEL 77.98 89.30 83.32 85.03 70.40 81.94 81.52 87.93 82.18
UniCorn 84.12 93.92 88.80 89.50 83.47 87.07 88.92 91.48 88.51
Table 13: Evaluation of text-to-image generation ability on GenEval benchmark. ‘Gen. Only’ stands for an image generation model, and ‘Unified’ denotes a model that has both understanding and generation capabilities. $\dagger$ refer to the methods using MLLM rewriter.The best Overall results are bolded.
[…]
Gen. Only PixArt-$\alpha$ (Chen et al., 2024b) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.12.1 (Rombach et al., 2022) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E $2$ (Ramesh et al., 2022) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu33-Gen  (Wang et al., 2024b) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL (Podell et al., 2024) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E $3$ (Betker et al., 2023b) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium (Esser et al., 2024) 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev (Labs, 2024) 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified Chameleon (Team, 2024) - - - - - - 0.39
LWM (Liu et al., 2024a) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X (Ge et al., 2024) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
TokenFlow-XL (Qu et al., 2024) 0.95 0.60 0.41 0.81 0.16 0.24 0.55
ILLUME (Wang et al., 2024a) 0.99 0.86 0.45 0.71 0.39 0.28 0.61
Janus (Wu et al., 2025b) 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Transfusion (Zhou et al., 2024a) - - - - - - 0.63
Emu33-Gen(Wang et al., 2024b) 0.99 0.81 0.42 0.80 0.49 0.45 0.66
Show-o (Xie et al., 2024) 0.98 0.80 0.66 0.84 0.31 0.50 0.68
Janus-Pro-7B (Chen et al., 2025d) 0.99 0.89 0.59 0.90 0.79 0.66 0.80
MetaQuery-XL (Pan et al., 2025) - - - - - - 0.80
BAGEL (Deng et al., 2025a) 0.99 0.95 0.76 0.87 0.50 0.60 0.78
Show-o2 (Xie et al., 2025c) 1.00 0.87 0.58 0.92 0.52 0.62 0.76
BAGEL (Deng et al., 2025a) 0.99 0.92 0.75 0.89 0.54 0.63 0.79
IRG*  (Huang et al., 2025c) 0.98 0.94 0.83 0.86 0.74 0.73 0.85
UniGen*  (Tian et al., 2025) 1.00 0.94 0.78 0.87 0.57 0.54 0.78
UniRL  (Mao et al., 2025) 0.96 0.80 0.67 0.86 0.50 0.67 0.74
UniCorn 0.99 0.94 0.80 0.88 0.61 0.73 0.82
GPT-4o (OpenAI, 2025) 0.99 0.92 0.85 0.92 0.75 0.61 0.84
Table 14: Quantitative evaluation results on DPG
[…]
PixArt-$\alpha$ (Chen et al., 2024b) 74.97 79.32 78.60 82.57 76.96 71.11
Lumina-Next (Zhuo et al., 2024) 82.82 88.65 86.44 80.53 81.82 74.63
Playground v2.5 (Li et al., 2024a) 83.06 82.59 81.20 84.08 83.50 75.47
Hunyuan-DiT (Li et al., 2024c) 84.59 80.59 88.01 74.36 86.41 78.87
Janus (Wu et al., 2025b) 82.33 87.38 87.70 85.46 86.41 79.68
Janus-Pro-1B (Chen et al., 2025e) 87.58 88.63 88.17 88.98 88.30 82.63
DALL-E 3 (Betker et al., 2023b) 90.97 89.61 88.39 90.58 89.83 83.50
FLUX.1-dev  (Labs, 2024) 74.35 90.00 88.96 90.87 88.33 83.84
SD3 Medium (Esser et al., 2024) 87.90 91.01 88.83 80.70 88.68 84.08
Janus-Pro-7B (Chen et al., 2025e) 86.90 88.90 89.40 89.32 89.48 84.19
BAGEL  (Deng et al., 2025a) - - - - - 84.03
UniCorn 91.62 91.97 91.39 91.22 91.64 86.83
Prompt for Proposer System Prompt: Character Introduction You are a specialist dataset architect for PromptBench. Your mission is to synthesize high-quality, high-complexity text-to-image prompts that push the limits of generative models. Your Task -Target Category: Generate prompt ONLY for the category defined by: {major category}. -Category Definition and Specific Rule(MUST FOLLOW THE RULE FOR THE TARGET CATEGORY): {category rule} -Informational Density: The prompt must contain sufficient descriptive detail to ensure complex image generation. Do not prioritize brevity over informational density Response Format Strictly follow the JSON format to output only the modified dialog without redundancy, and do not add comments (//) in the response. { "major_category": "The primary classification", "subcategory": "The secondary classification" "prompt": "The high-density descriptive instruction." } Example {Few-shot Example} User Prompt: Generate exactly ONE new prompt. Target Major Category: {major category}. Target Subcategory: {subcategory} Each generated item must have a major_category field set to {major category}, a subcategory field set to {subcategory}, and a prompt field. Ensure high diversity and strictly adhere to the rule.
Figure 8: The prompt template for prompt proposer.
Prompt for Image Judge System Role: You are a rigorous Visual Quality Assessment Expert. Your mission is to evaluate the alignment and technical fidelity of generated images against specific text prompts using a deterministic, objective framework. Evaluation Criteria (Ranked by Priority): {category specific Judgement Rubrics } Scoring Standard:(0 - 10) {category specific scoring standard} Response Format: Return a strictly valid JSON object only. Do not include conversational filler, markdown commentary, or code block delimiters. { "analysis": "A concise, objective breakdown of the evaluation points.", "score": "Integer or float from 0 to 10" } Input Data: Category: {major category} Prompt: {prompt} Image: [Image]
Figure 9: The prompt template for reward judger.
Method External Model Free External Data Free External Model Hyperparameters↓
IRG GPT-4o+Qwen2.5VL 0
UniRL GPT-4o 1
SRUM SAM3 1
RecA GPT-4o 3
UniCorn - 0
Table 15: Comparison of different methods in terms of external dependencies and prompt construction strategies. Without relying on external task-specific models or annotated data, UniCorn achieves state-of-the-art performance on OneIG-EN using only 5K training samples.
Figure 10: Cases of UniCycle.
Prompt for UniCycle evaluation (non-text task type) You are a strict visual QA evaluation assistant. You will be given: 1) TASK_TYPE: the evaluation dimension to consider. 2) IMAGE_PROMPT describing what the image should contain. 3) ONE QA pair (Question, Answer). 4) A Reference Answer. Your Task
Determine whether the Answer is consistent with IMAGE_PROMPT for TASK_TYPE only.
Ignore all other aspects. You may use the Reference Answer only for equivalence checking. Rules
- Use ONLY IMAGE_PROMPT; do NOT use external knowledge.
- Output "yes" ONLY if IMAGE_PROMPT clearly supports the Answer for TASK_TYPE. - Output "no" if the Answer contradicts IMAGE_PROMPT, or if IMAGE_PROMPT is insufficient. - Output "no" if the Answer is a refusal, uncertainty, or hedging. - Be strict: required details must be explicitly supported. - Do NOT explain. Output JSON only. Normalization rules (for equivalence checking only)
- Ignore letter case, punctuation, and extra whitespace.
- Minor spelling variants are equivalent (e.g., gray/grey, color/colour). Output JSON with exactly these keys: { "question": "<question>", "answer": "<answer>", "evaluation": "yes" or "no" } [TASK_TYPE] {task_type} [IMAGE_PROMPT] {image_prompt} Question: {question} Answer: {answer} Reference Answer: {refer_ans}
Figure 11: The prompt template for UniCycle evaluation(non-text task type).
Prompt for UniCycle evaluation (text task type) You are a strict text rendering QA evaluator. You will be given:
1) IMAGE_PROMPT describing what the image should contain
2) ONE QA pair (Question, Answer) 3) A Reference Answer Your task:
Count how many required words in the Answer are correctly supported by IMAGE_PROMPT and Reference Answer.
Use ONLY IMAGE_PROMPT. Do NOT use external knowledge. Output JSON only with exactly these keys: { "question": "<question>", "answer": "<answer>", "evaluation": "<number of correctly answered words>" } [IMAGE_PROMPT] {image_prompt} Question: {question} Answer: {answer} Reference Answer: {refer_ans}
Figure 12: The prompt template for UniCycle evaluation (text task type).
Question Type Count Ratio (%)
Total questions 2968 100.00
MCQ questions 1067 35.95
Yes/No questions 200 6.74
Open-ended questions 1701 57.31
Table 16: Question types distribution of UniCycle.
Model Bagel Show-o2 Janus-Pro UniCorn* UniCorn
Soft score 58.2 52.5 25.8 58.6 66.6
Table 17: Soft score results on UniCycle.
Figure 13: Failure cases of UniCorn in chanllenging tasks of Negation and Counting.